OpenStudy (xenesis98):

Suppose two drugs are routinely used for treatment of a particular kidney disorder. Drug 1 is known to cure the disease 85% of the time and costs $90. Drug 2 is known to cure the disease 70% of the time and costs $65. The two drugs work independent of each other (that is, administration of one has no effect on the efficacy of the other). The two treatment plans are as follows: Plan A: Treatment with Drug 1—if not effective, treatment with Drug 2. Plan B: Treatment with Drug 2—if not effective, treatment with Drug 1. Which statement is most correct in this situation?

OpenStudy (xenesis98):

a.Based on the overall probability of a cure, Plan A should be selected over Plan B. b.Based on the overall probability of a cure, Plan B should be selected over Plan A. c.Based on the overall cost of treatment, Plan A should be selected over Plan B. d.Based on the overall cost of treatment, Plan B should be selected over Plan A. e.Based on the probability of a cure and the cost of treatment, both plans are equivalent, so either can be selected.

OpenStudy (kropot72):

The overall probability of a cure with plan A is given by: \[\large P(cure\ with\ plan\ A)=0.85+0.15\times0.7=0.955\] The overall probability of a cure with plan B is given by: \[\large P(cure\ with\ plan\ B)=0.7+0.3\times0.85=0.955\] Therefore both plans have the same probability of a cure. Let the cost of a cure be given by the variable X. With plan A: \[\large E(X)=90\times0.85+(90+65)\times(0.15\times0.7)=$92.78\] With plan B: \[\large E(X)=65\times0.7+(90+65)\times(0.3\times0.85)=$85.03\] Can you now make the correct choice of answer?

OpenStudy (xenesis98):

D?

OpenStudy (kropot72):

Yes, D is correct.